Pressure gauge indicate that the pressure at point 1 is 10kgPa greater than the pressure at point 2. The cross sectional area at point 2 is 1/5th the cross sectional area at point 1. Determine the speed of the air at point 2
1
Expert's answer
2020-12-14T07:31:32-0500

P1 - is pressure which pressure gauge indicates at the point 1

P2 - is pressure which pressure gauge indicates at the point 2

"\Delta"P = 10kP

P1 = P2 - "\Delta"P

S2 = S1 / 5; S1 = 5S2

T = const in pressure gauge

according to ideal gas law

PV = const

P1S1H = P2V2H -> (P2 - "\Delta"P) *5S2 = P2 S2

P2 = 1.25P1 = 12500Pa

v = "\sqrt{3P/\rho}" = >

"\rho" = 1.29kg/m3 - for air

v = "\sqrt{3 * 12500/1.29}" = 170m/s